The velocity of a tsunami in feet per period


6


9


12


15


18


21


24



Height of tsunami in feet


4


9


16


25


36


49


64



Use the data in the above in inductive reasoning to answer each of the following questions



1. If a pendulum has a length of 49 units what is its period?


2. If the length of a pendulum is quadrupled what happens to its period?



Verify that each of the following statements is a false statement by finding a counterexample



1. X/y=1


2. X+y/ 3= x+1


3. 3 √x2+16= x+4






1
Expert's answer
2021-11-22T13:54:27-0500

1.1.


"2=\dfrac{12}{6}=>\dfrac{16}{4}=4"

"2=\dfrac{18}{9}=>\dfrac{36}{9}=4"

"2=\dfrac{24}{12}=>\dfrac{64}{16}=4"

Height of a tsunami increases by 4 times when the velocity is doubled. 


1.2.


"2=\dfrac{30}{15}=>\dfrac{h}{25}=4=>h=100"

The height of a tsunami should be 100 ft if its velocity is 30 feet per second.


2. The period of a simple pendulum "T" is 


"T=2\pi \sqrt{L/g},"

where "L" is the length of the string and "g" is the acceleration due to gravity.

2.1. Given "L=49\ m, g=9.8\ m/s^2"


"T=2\pi\sqrt{49\ m/(9.8\ m/s^2)}=2\sqrt{5}\pi
 \ s\approx14.050\ s"

2.2.


"T=2\pi \sqrt{L/g}"

If you quadruple the length of the string,  the period doubles.


3.1.

Let "x=2, y=1." Then


"x/y=2/1=2\not=1"

The given statement is False.


3.2.

Let "x=1, y=6." Then


"x+y\/3=1+6\/3=3"




"x+1=1+1=2"

"x+y/3=3\not=2=x+1"

The given statement is False.


3.3.

Let "x=0." Then


"3\sqrt{x^2+16}=3\sqrt{(0)^2+16}=12"




"x+4=0+4=4"

"3\sqrt{x^2+16}=12\not=4=x+4"

The given statement is False.